40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5453     Entry time: Mon Sep 19 01:38:00 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: today's locking activity 

[Anamaria / Kiwamu]

 The incident beam pointing was improved by using PZT1 and PZT2. 

With some triggers the lock of PRMI became smoother.

For the DRMI lock, the MICH and SRCL signals on AS55 are not quite decoupled, so we should find cleaner signals for them.

 

(what we did)

 + locked the Y arm

 + aligned incident beam by using PZT1(#5450) and PZT2. The spot positions on ITMY and ETMY are now well-centered.

 + tried activating C1ASS but failed. It needs some gain changes due to the new PZT1 response.

 + locked the X arm

 + aligned the TRX PD (Thorlab signal PD) and set the trigger.

 + C1ASS also doesn't work for the X arm

 + realigned the PRM and BS oplevs. the PRM oplev was clipped at a steering mirror on the optical bench

 + locked PRMI and aligned the PRM mirror such that the optical gain was maximized

 + optimized the demod phase of AS55 and REFL11

 + checked the UGF of the MICH and PRCL lock. The UGF of MICH is about 100Hz with gain of -20, and the UGF of PRCL was 85 Hz with gain of 0.1

 + adjusted the output matrix such that the MICH control doesn't couple into the PRCL control.

 + set the triggers for the MICH and PRCL control to make the lock acquisition smoother.

 + tried locking DRMI and it was sort of locked. However the SRCL signal showed up a lot in AS55_Q, where the MICH signal is extracted.

ELOG V3.1.3-